# Data Readiness for AI: A 360-Degree Survey

### v1

@@ Abstract.
Data are the critical fuel for Artificial Intelligence (AI) models. Poor quality data produces inaccurate and ineffective AI models that may lead to incorrect or unsafe use. Checking for data readiness is a crucial step in improving data quality. Numerous R&D efforts have been spent on improving data quality. However, standardized metrics for evaluating data readiness for use in AI training are still evolving. In this study, we perform a comprehensive survey of metrics used for verifying AI’s data readiness. This survey examines more than 120 papers that are published by ACM Digital Library, IEEE Xplore, other reputable journals, and articles published on the web by prominent AI experts. This survey aims to propose a taxonomy of data readiness for AI (DRAI) metrics for structured and unstructured datasets. We anticipate that this taxonomy can lead to new standards for DRAI metrics that would be used for enhancing the quality and accuracy of AI training and inference.
Data readiness for AI, Data quality metrics, Data quality assessment
Authors’addresses: K. Hiniduma and S. Byna, The Ohio State University, Columbus, OH; emails: hiniduma.1@osu.edu, byna.1@osu.edu; J. L. Bez, Lawrence Berkeley National Laboratory, Berkeley, CA; email: jlbez@lbl.gov
copyright: acmcopyrightjournalyear: 2022doi: XXXXXXX.XXXXXXXconference: ACM Computing Surveys; June 03–05, 2022; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Machine learningccs: Social and professional topics Quality assurance
@@ 1. Introduction
Data readiness for artificial intelligence (AI) refers to the critical process of preparing and ensuring the quality, accessibility, and suitability of datasets before using them for artificial intelligence applications. Readying the data is a fundamental step, which involves collecting, cleaning, organizing, and validating the dataset not only to make them compatible with AI algorithms and models but also to make certain that the datasets are appropriate and unbiased. By achieving data readiness, organizations can maximize the accuracy, efficiency, and effectiveness of their AI systems, ultimately leading to more informed decision-making and successful AI-driven outcomes.
@@
Data readiness for AI is an important concern in the field of AI applications, as evidenced by a survey conducted by Scale AI (Hwang, 2022). The survey highlights that a significant number of participants encountered challenges related to data readiness within their machine learning (ML) projects. Similarly, a comprehensive study (Woodie, 2020) involving nearly 2,400 respondents from over 100 countries explains the time-intensive nature of data preparation and cleaning tasks for data scientists working with AI applications. Moreover, this study reveals that these activities toward readying data can adversely affect job satisfaction. It is crucial to recognize that the quality of outcomes generated by an AI system is strongly linked to the readiness of the input data. This connection highlights the significance of addressing the ”garbage in, garbage out” saying, which emphasizes that flawed or insufficient input data will inevitably lead to inaccurate and unreliable results from AI algorithms (Schmelzer, 2019). Hence, ensuring the availability of high-quality and well-prepared data for training ML models is crucial, as it leads to more precise and dependable predictions.
@@
With growing requirements of unbiased high-quality data for AI, the field of quantitative evaluation of data quality with appropriate metrics is still evolving. IBM has developed the Data Quality Toolkit (Developer, 2021) provides a suite of tools and functionalities to streamline the preparation and cleaning of their data. A notable feature of the IBM Data Quality Toolkit is the Data Readiness Report, which is generated automatically after the data preparation and cleansing process. This report offers insightful observations into the quality and readiness of the data for use in AI applications, including data quality metrics defined by Sidi et al. (Sidi et al., 2012) such as completeness, consistency, accuracy, and timeliness, along with suggestions for improvement. On the other hand, Ravi et al. (Ravi et al., 2022) focus on the critical process of making experimental datasets FAIR (Findable, Accessible, Interoperable, Reusable) for AI readiness. Throughout their analysis, they emphasize the usage of current data infrastructure to establish a framework suitable for automatic AI-powered exploration. To achieve this objective, they publish FAIR and AI-ready datasets. The published datasets are at The Materials Data Facility (Blaiszik et al., 2016). Additionally, the study complements these efforts by illustrating how FAIR and AI-ready datasets can be effectively used for AI inference.
@@
Although separate efforts and tools mentioned above are available to prepare data for AI, there is a lack of a comprehensive study on effective metrics for various types of data. To address that gap, we perform a comprehensive examination of the existing data readiness metrics for AI, covering both structured and unstructured data dimensions, and discuss on the existing tools and techniques employed to assess data readiness. Furthermore, we explore in this survey the metrics designed to evaluate fairness and privacy related issues in data that could impact the decision-making processes of AI algorithms. The significance of not incorporating social-minded measures such as fairness and bias into data quality processes has the potential to compromise the integrity of the data quality.
Figure 1. 360° View of Mapping Data Readiness Dimensions for AI
Our survey draws upon a comprehensive set of data quality dimensions and data preparation techniques by specifically focusing on defining the primary data readiness dimensions essential for AI applications. We aim to identify and showcase the existing metrics and scoring mechanisms available in the literature that effectively measure the readiness of data for AI. By thoroughly exploring these dimensions and metrics, our survey provides valuable insights into assessing the preparedness of data for AI, contributing to a comprehensive understanding of data readiness. Fig.  1 illustrates the 360° plot of data readiness dimensions for AI, encompassing data evaluation metrics discussed in this study.
@@
Given the critical importance of data readiness for AI metrics, it is essential to stay up to date with the most effective research and advancements in this area. We have performed an extensive search for available data readiness metrics by sifting through the literature available in ACM Digital Library, IEEE Xplore, and other reputable journals in computer science. Considering that the field of assessing data readiness is still nascent, we have also surveyed several web articles that are available online. Overall, we have surveyed nearly 30 papers from ACM Digital Library, over 20 papers from IEEE Xplore, 10 papers each from Springer and Science Direct, and more than 40 papers from other reputable journals worldwide. Additionally, we highlight discussions on six web articles and explore the metrics used in four commercially used tools in this field. In Table 1, we present the search queries used to procure the sources for this study, categorized into general, structured data, and unstructured data search queries.
In shaping this study, our focus encompassed sources published across different timeframes – pre-2000, 2000-2010, and post-2010. Fig. 2 illustrates the distribution of sources across different timeframes. Paying attention to references prior to 2000 is particularly important because they include some well-established data readiness metrics that have stood the test of time and remain relevant today. These earlier references help provide context and insights into the evolution of data readiness metrics over the years.
Figure 2. Illustration of Source Distribution Across Timeframes
@@
”data readiness” AND ”AI” ”data quality” AND ”assessment”
@@
e.g., ”discriminat*” AND ”metric” OR ”measure” OR ”evaluat*”
”speech quality” AND ”metric” OR ”measure” OR ”evaluat*”
@@
”data readiness” ”data quality dimension” ”audio quality” AND ”metric”OR ”measure” OR ”evaluat*”
@@
”data readiness” AND ”machine learning” OR ”ML” ”data quality” AND ”metric” ”video quality” AND ”metric”OR ”measure” OR ”evaluat*”
@@
”AI ready” ”data prepare” AND ”AI” ”image quality” AND ”metric”OR ”measure” OR ”evaluat*”
@@
”data quality” AND ”machine learning” ”data read*” AND ”metric” ”visual quality” AND ”metric” OR ”measure” OR ”evaluat*”
”data quality AND ”measure” ”data preprocess”
”data quality” AND ”evaluation” ”data clean”
”data quality” and ”AI” ”data quality” AND ”survey”
In the remainder of the paper, we describe studies that target understanding metrics for AI data readiness. We provide our definition of data readiness for AI and how it can be used to shape AI pipelines. In Section 4, we introduce a comprehensive set of metrics being used for measuring the readiness of structured and unstructured data. In Section 5, we list a set of gaps and challenges in defining data readiness for AI and then conclude the discussion with future work.
2. Related Work
@@
Toward gaining an understanding of metrics for AI data readiness, a small number of studies surveyed data quality and preprocessing stages in preparing data for AI. Priestley et al. (Priestley et al., 2023) conducted a significant study highlighting the critical role of decision-makers and practitioners in driving the shift toward data-focused practices. They emphasized the need for high-quality data and underscored the significance of data cleaning and preprocessing stages, including feature selection, duplicate elimination, outlier removal, consistency assurance, and handling of missing values. Documentation of these preprocessing steps was deemed essential to ensure compatibility and identify potential dependencies and information leakages among features. Their insights, derived from an extensive literature survey, laid the foundation for recognizing the pivotal role of data quality in AI applications. Yalaoui et al. (Yalaoui and Boukhedouma, 2021) presented a survey paper specifically focused on data quality evaluation and enhancement. They compared existing frameworks, models, and methods for data quality evaluation and enhancement, and identified research directions and challenges.
@@
There are a few studies that looked into metrics for unstructured data. Elmagarmid et al. (Elmagarmid et al., 2007) explored into similarity metrics for duplicate record detection, specifically focusing on challenges posed by typographical variations in string data. Li et al. (Li et al., 2017) concentrated on feature selection metrics, categorizing traditional approaches into wrapper methods, filter methods, embedded methods, and hybrid methods. Furthermore, Forman (Forman, 2003) explored feature selection metrics for text classification, examining their performance. The studies on image quality measures (Thung and Raveendran, 2009) and perceptual visual quality metrics (Lin and Jay Kuo, 2011) play a crucial role in understanding data readiness for unstructured data, particularly in the context of visual data, where image quality assessment and perceptual metrics are essential for effective AI applications.
@@
In the context of bias and fairness, Shahbazi et al. (Shahbazi et al., 2023) provided a survey of techniques focused on identifying and mitigating representation bias across diverse data types, such as structured data, image data, textual data, and graph data. They discussed the problem definition, causes, and methods for measuring and quantifying representation bias in both structured and unstructured data. Addressing the issue of discrimination in data readiness for AI systems, Ntoutsi et al. (Ntoutsi et al., 2020) focused on the challenges and implications of biased data on the fairness and accuracy of AI-based decision-making. They emphasized the importance of understanding and mitigating biases in data to prevent the perpetuation of discriminatory practices. Similarly, in the context of privacy, Wagner et al. (Wagner and Eckhoff, 2018) reviewed over 80 privacy metrics across six domains categorizing them based on the aspect of privacy measured, required inputs, and data types needing protection. It identifies research gaps in areas such as metric combination and interdependent privacy, proposing a method for selecting appropriate metrics through nine key questions. Emphasizing the importance of employing multiple metrics to address diverse privacy facets, the paper serves as a reference guide and toolbox for privacy researchers, aiding informed choices in metric selection for specific scenarios.
@@
Building upon existing comprehensive surveys, our study aims to contribute to the evolving field of data readiness metrics, with a specific focus on metrics for both structured and unstructured data. While previous works have explored specific dimensions or focused on individual metrics, we address the lack of a comprehensive study including numerous data readiness dimensions for AI applications. There are broader perspectives in addressing data readiness for AI applications. We identify numerous crucial factors that define data readiness, such as data preparation, privacy leakage evaluation, data discrimination evaluation, compliance to FAIR principles, mislabeled data detection, feature relevancy analysis, bias-related issues, and quality evaluation of speech and multimedia data. To survey existing literature in these dimensions and to identify gaps, our effort aims to advance the AI data readiness field and inform best practices for evaluating the suitability of data for AI applications. Given the growing importance of data readiness, our study fills a crucial gap in the literature and provides valuable insights for practitioners and decision-makers in the field.
3. Data Readiness for AI (DRAI)
In this section, we examine the concept of data readiness for AI and explore the processes involved in preparing data for effective AI usage. Furthermore, we will discuss the significance of assessing data readiness for AI, emphasizing the importance of robust frameworks, methodologies, and metrics to evaluate and enhance the suitability of data for AI applications. By addressing the key challenges and considerations in data readiness, organizations can lay a solid foundation for AI implementation and unlock the full potential of their data resources.
3.1. Processes of Readying Data for AI
The AI community defines the processes of readying data for AI as a series of steps aimed at preparing the data to be suitable for machine learning and AI applications. According to IBM Biddle et al. (Biddle and Christensen, nd) describe that data preparation for AI involves transforming data into a format suitable for model development, measurement, and training. This includes processes such as selecting relevant attributes, merging data sets, deriving new attributes, formatting, sorting the data, addressing missing values, and ensuring data accuracy and completeness. In parallel, Informatica (Tawfik et al., 2021) emphasizes critical processes in data preparation for AI projects. These processes include finding and accessing trusted data, analyzing data lineage and relationships, preparing and cleansing the data, enriching the data through feature engineering or incorporating additional data, and operationalizing the data pipeline for reuse and automation. Similarly, Lawton (Lawton, 2022) characterizes data preparation as a time-consuming and complex process. It involves various processes such as problem formulation, data collection, data exploration, data cleansing and validation, data structuring, feature engineering, and feature selection. These processes collectively aim to ensure accurate and adaptable algorithms, ultimately saving time and effort for data scientists and business users.
To further enhance the quality, relevance, and usability of data for AI and machine learning applications, AltexSoft (Editor, 2020) outlines ten fundamental processes for preparing datasets. These processes incorporate problem articulation, establishing data collection mechanisms, assessing data quality, formatting data for consistency, reducing data, data cleaning, creating features, joining transactional and attribute data, rescaling data, and discretizing data.
In addition to these essential processes, the Data Readiness Report, as introduced by Afzal et al. (Afzal et al., 2020), assumes a crucial role in readying data for AI. This report serves as documentation of operations conducted during data preparation, cleaning, and quality analysis. By providing insights into dataset characteristics and evaluating them against relevant quality dimensions, the Data Readiness Report effectively addresses upfront data challenges and contributes to the overall effectiveness of data preparation for AI applications.
Overall, the AI community defines the processes of readying data for AI as a set of tasks that involve transforming, cleaning, validating, enriching, and documenting the data to make it suitable for model development, training, and analysis. These processes aim to ensure data quality, accuracy, relevance, and consistency, ultimately facilitating effective AI and machine learning outcomes.
3.2. Assessing Data Readiness
Assessing data readiness for AI is crucial for ensuring the performance, fairness, and reliability of AI systems. It involves evaluating quality, completeness, representativeness, and other data readiness dimensions, including feature relevancy, duplicity, class separability, FAIR (Findable, Accessible, Interoperable, and Reusable) compliance, and outliers used for training and deploying AI models. By conducting a thorough assessment, AI users can identify and address data quality issues, mitigate bias and fairness concerns, improve model generalization and robustness, and comply with data governance and ethical guidelines. In addition, assessing data readiness in the early stages of the ML pipeline is highly beneficial. By evaluating data readiness at an early stage, organizations can identify any data-related challenges or limitations upfront, allowing for proactive measures and adjustments. In Fig.  3, we illustrate the placement of data readiness evaluation within the ML pipeline, emphasizing the importance of conducting these assessments early on. Ultimately, assessing data readiness enhances the overall effectiveness and trustworthiness of AI systems, enabling them to make accurate predictions and informed decisions in real-world applications.
Figure 3. Enhanced ML pipeline with the evaluation of data readiness
Assessing data readiness for AI applications is a multifaceted challenge, as discussed by Allen (Allen, 2019) and Gupta et al. (Gupta et al., 2021). These complementary approaches provide insights into various aspects of data readiness. On the one hand, Allen’s article highlights five key dimensions: data accessibility, size, usability, understandability, and maintainability. It underscores the need for strategic data management, including external sources and adherence to good data practices. On the other hand, Gupta et al.’s toolkit offers a practical, automated approach to assessing data quality and addressing data readiness challenges. It provides an array of quality metrics for structured datasets, covering dimensions like completeness, accuracy, consistency, and timeliness. These metrics empower data scientists to understand data limitations and guide data preparation effectively. By merging these studies, there is potential for a comprehensive framework that improves data readiness in AI projects, including both strategic considerations and practical data quality assessment techniques.
Data scientists often face challenges during the pre-processing stage, which can significantly impact productivity and introduce data quality issues. These include dealing with missing or incomplete data, handling noisy or irrelevant features, addressing class imbalances in the dataset, ensuring data privacy and security, managing data scalability and storage constraints, handling unstructured or semi-structured data formats, and dealing with inconsistencies or errors in data sources. Afzal et al. (Afzal et al., 2020) introduced the concept of a data readiness report seeking to address these challenges. This report provides detailed insights into the quality of input data across various dimensions. It captures data properties, identifies quality issues, and documents data operations performed by different individuals involved in the data preparation process. Alongside Datasheets (Gebru et al. (Gebru et al., 2018)) and Model Cards (Mitchell et al. (Mitchell et al., 2019)), the data readiness report enhances transparency, documentation, and governance in the AI pipeline. It promotes responsible AI development by ensuring stakeholders have a clear understanding of the dataset’s strengths, limitations, and suitability for training machine learning models.
To further enhance the assessment of data readiness for AI, Lawrence (Lawrence, 2017) focuses on assessing data readiness by categorizing it into different bands (C, B, and A) based on accessibility, accuracy, representation, and appropriateness. Each band represents a specific level of data quality and readiness for analysis or distribution. Band C focuses on data accessibility, requiring it to be made machine-readable and ethically handled. Band B emphasizes verifying data accuracy and representation, addressing missing values and errors. Band A considers the appropriateness of the dataset for specific tasks involving careful evaluation and potential annotation or new data collection. By categorizing data readiness into these bands, the framework provides a structured methodology for evaluating and improving data quality, thereby enhancing the reliability and effectiveness of AI systems.
Similar to Lawrence’s (Lawrence, 2017) work on data readiness, Harvey et al. (Harvey and Glocker, 2019) also addressed the topic by introducing the MIDaR (Medical Imaging Data Readiness) scale. It is a framework for assessing the readiness of medical imaging data for AI applications, considering FAIR (Findable, Accessible, Interoperable, and Reusable) principles and different stages of data preparation. It categorizes data into four levels: D, C, B, and A. Level D represents abundant but inaccessible and un-anonymized data. Level C indicates anonymized data with artifacts and noise. Level B signifies structured and quality-controlled data without labels, and Level A represents labeled and statistically powered data ready for AI tasks. As data progresses toward Level A, its value increases, although the volume may decrease. The MIDaR scale provides insights into data quality, ethics, labeling, and statistical power when preparing medical imaging data for AI analysis.
In line with the focus on data readiness assessment, Patel et al. (Patel et al., 2022) (Patel et al., 2023) underscore the importance of systematically analyzing data quality to improve the efficacy of machine learning and artificial intelligence applications. Their research emphasizes the need for intelligently designed metrics and transformation operations that effectively address data quality gaps, reducing the effort required for debugging in the machine learning pipeline. This perspective on data quality aligns with the principles embedded in OpenDataVal introduced by Jiang et al. (Jiang et al., 2023), an existing benchmark framework. OpenDataVal provides a unified environment for researchers and practitioners to apply and compare various data valuation algorithms. With a diverse collection of datasets, the implementation of state-of-the-art data valuation algorithms, and proposed machine learning tasks for evaluating data quality, OpenDataVal significantly contributes to streamlining data preparation. The framework’s integrated environment and prediction model API further facilitate the seamless integration of various models. This comprehensive approach in OpenDataVal aims to enhance machine learning outcomes by promoting robust data preparation and evaluation strategies.
Table 2. Comparison of Data Readiness Assessment Approaches for AI Applications
Reference Summary
Allen (Allen, 2019)
Explores challenges and considerations in assessing data readiness for AI applications.
Focuses on data accessibility, size, usability, understandability, and maintainability.
Emphasizes the importance of data strategy, considering external sources, and following good data practices.
Gupta et al.(Gupta et al., 2021)
Presents a data quality toolkit for automated data quality assessment.
Reduces preparation time and enhances data quality.
Offers various quality metrics for accessible and actionable assessment by diverse users.
Assists in identifying issues and guiding data cleansing and transformation.
Afzal et al. (Afzal et al., 2020)
Introduces the concept of a data readiness report.
Provides comprehensive documentation of data quality across dimensions.
Captures data properties, quality issues, and operations during data preparation.
Enhances transparency, documentation, and governance in the AI pipeline.
Lawrence (Lawrence, 2017)
Categorizes data readiness into bands (C, B, and A) based on accessibility, accuracy, representation, and appropriateness.
Each band represents a specific level of data quality and readiness for analysis or distribution.
Harvey et al. (Harvey and Glocker, 2019)
Introduces the MIDaR scale for assessing medical imaging data readiness for AI applications.
Considers FAIR principles and data preparation stages.
Categorizes data into four levels (D, C, B, and A) based on data quality, ethics, labeling, and statistical power.
Patel et al. (Patel et al., 2022) (Patel et al., 2023)
Highlights the importance of systematically analyzing data quality to improve AI application efficacy.
Provides intelligently designed metrics and transformation operations for addressing data quality gaps.
Reduces debugging effort in the machine learning pipeline.
Jiang et al. (Jiang et al., 2023)
OpenDataVal creates a unified framework where practitioners can employ and compare diverse data valuation algorithms
Offer a broad range of datasets, implementing data valuation algorithms, and defining machine learning tasks for assessing data quality.
In Table 2, we compare various AI readiness approaches discussed in this section. Evaluating data readiness for AI is crucial for robust and reliable machine learning systems. This process – includind data quality, completeness, representativeness, and more – aids in addressing issues, reducing biases, enhancing model performance, and ensuring ethical compliance. The complexities involved emphasize the need for a strong data strategy, exploration of external data sources, and robust data pipelines. The advent of assessment frameworks and toolkits empowers practitioners to improve the trustworthiness of AI systems. Ongoing research in data readiness will contribute to responsible AI deployment, enabling accurate decisions across diverse applications. Prioritizing data readiness establishes a solid foundation for AI success, promoting transparency, accountability, and responsible development.

### v2

@@ Abstract.
Artificial Intelligence (AI) applications critically depend on data. Poor quality data produces inaccurate and ineffective AI models that may lead to incorrect or unsafe use. Evaluation of data readiness is a crucial step in improving the quality and appropriateness of data usage for AI. R&D efforts have been spent on improving data quality. However, standardized metrics for evaluating data readiness for use in AI training are still evolving. In this study, we perform a comprehensive survey of metrics used to verify data readiness for AI training. This survey examines more than $140$ papers published by ACM Digital Library, IEEE Xplore, journals such as Nature, Springer, and Science Direct, and online articles published by prominent AI experts. This survey aims to propose a taxonomy of data readiness for AI (DRAI) metrics for structured and unstructured datasets. We anticipate that this taxonomy will lead to new standards for DRAI metrics that would be used for enhancing the quality, accuracy, and fairness of AI training and inference.
Data readiness, Data quality metrics, AI-ready data
copyright: acmcopyrightjournalyear: 2022doi: XXXXXXX.XXXXXXXconference: ACM Computing Surveys; June 03–05, 2022; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Data readiness for artificial intelligence (AI) refers to the critical process of preparing and ensuring the quality, accessibility, and suitability of datasets before using them for AI applications. Readying the data is a fundamental step, which involves collecting, cleaning, organizing, and validating the dataset not only to make them compatible with AI algorithms and models, but also to make certain that the datasets are appropriate and unbiased. By achieving data readiness, organizations can maximize the accuracy, efficiency, and effectiveness of their AI systems, ultimately leading to more informed decision-making and successful AI-driven outcomes.
@@
Data readiness for AI (DRAI) is an important concern in AI applications, as evidenced by a survey conducted by Scale AI (Hwang, 2022). A significant number of participants encountered challenges related to data readiness within their machine learning (ML) projects. Similarly, a study (Woodie, 2020) involving nearly $2,400$ respondents from over $100$ countries explains the time-intensive nature of data preparation for data scientists working with AI applications. It is crucial to recognize that the quality of outcomes generated by an AI system is heavily linked to the readiness of the input data. This connection highlights the significance of addressing the “garbage in, garbage out” saying, which emphasizes that flawed or insufficient input data will inevitably lead to inaccurate and unreliable results from AI algorithms (Schmelzer, 2019). Hence, ensuring the availability of well-prepared data for training machine learning (ML) models is critical, as it leads to more precise and dependable predictions.
@@
With growing requirements of unbiased data for AI, the field of quantitative evaluation of data readiness with appropriate metrics is still evolving. Data Quality Toolkit (DQT) (Developer, 2021) provides a suite of tools and functionalities to streamline the preparation and cleaning of data. DQT’s data quality report various dimensions of data quality metrics defined by Sidi et al. (Sidi et al., 2012) such as completeness, consistency, accuracy, and timeliness, along with suggestions for improvement. Ravi et al. (et al., 2022) focus on the critical process of making experimental datasets FAIR (Findable, Accessible, Interoperable, Reusable) for AI readiness. They emphasize the usage of current data infrastructure to establish a framework suitable for automatic AI-powered exploration. To achieve this objective, they publish FAIR and AI-ready datasets (Blaiszik et al., 2016). This study also illustrates usage of FAIR principle compliance (FAI, 2024) and AI-ready datasets for inference.
@@
Although these separate efforts and tools are available to improve quality of data, there is a lack of a comprehensive study on effective metrics and standards for evaluating data readiness for AI. To address that gap, we perform a comprehensive examination of the existing metrics and tools that could be used for evaluating data readiness, covering both structured and unstructured data dimensions We also describe the metrics designed to evaluate fairness and privacy related issues in data, which critically impact the process of decision-making in AI algorithms.
This survey refers to a comprehensive set of data readiness dimensions and data preparation techniques targeting data usage in AI. Metrics targeting data readiness for AI (DRAI) contain a subset of data quality dimensions including completeness, duplicates, correctness, and timeliness. This distinction between DRAI and data quality is critical for understanding our survey’s scope. We identify the existing metrics and scoring mechanisms (§3) and existing frameworks or tools (§4) in the literature that could be used to measure DRAI. Based on the distillation of available literature, we propose a potential comprehensive definition of data readiness for AI using six dimensions (§5). We discuss gaps and challenges towards developing a DRAI assessment framework (§6). This survey is particularly aimed at data preparers for future AI use and data scientists who analyze the data to ensure their datasets are ready for AI applications.
2. Scope of the Study
@@
Our literature review methodology was carefully structured to ensure a comprehensive and unbiased examination of existing research. The search queries used to obtain the sources for this study are presented in Table 1, categorized into general, structured data, and unstructured data search queries. As a result, we gathered nearly $30$ papers from ACM Digital Library, over $20$ papers from IEEE Xplore, $10$ papers each from Springer and Science Direct, and more than $40$ papers from journals including Nature, Springer and Science Direct, as well as several relevant books, to review. Additionally, we highlight discussions on six web articles and explore the metrics used in six commercially used tools.
@@
“data readiness” AND “AI” “data quality” AND “assessment”
@@
e.g., “discriminat*” AND “metric” OR “measure” OR “evaluat*”
“speech quality” AND “metric” OR “measure” OR “evaluat*”
@@
“data readiness” “data quality dimension” “audio quality” AND “metric” OR “measure” OR “evaluat*”
@@
“data readiness” AND “machine learning” OR “ML” “data quality” AND “metric” “video quality” AND “metric” OR “measure” OR “evaluat*”
@@
”AI ready” “data prepare” AND “AI” “image quality” AND “metric” OR “measure” OR “evaluat*”
@@
“data quality” AND “machine learning” “data read*” AND “metric” “visual quality” AND “metric” OR “measure” OR “evaluat*”
“data quality” AND “measure” “data preprocess”
“data quality” AND “evaluation” “data clean”
“data quality” and “AI” “data quality” AND “survey”
Papers and articles were included if they discussed data readiness metrics for AI or data quality metrics, covered structured or unstructured data dimensions, and addressed fairness and privacy issues related to DRAI. We included sources that provided metrics or tools for evaluating data readiness and those that were peer-reviewed or published in reputable journals. We also used web articles that provided valuable insights and were considered highly relevant to our study. Additionally, if the metrics were related to AI data preprocessing, such as feature relevancy, class imbalance measures, and FAIR compliance, they were also included.
We excluded papers that did not focus on DRAI, those that were not peer-reviewed, and articles that lacked substantial contributions to understanding DRAI metrics. We also excluded duplicates and studies not available in English. Additionally, we excluded papers if the metrics were not quantifiable or unrelated to the pre-data training stage in AI.
Our survey aims to identify and analyze the existing metrics and scoring mechanisms for measuring DRAI, covering both conventional dimensions of data quality (e.g., completeness, outliers, timeliness, and correctness) and AI-specific dimensions (e.g., fairness, feature importance, class imbalance, and mislabeled data). We seek to provide insights into how these dimensions and metrics can be used to assess the preparedness of data for AI applications.
Figure 1. Papers chosen for this survey from different time frames.
In forming this study, our focus included published literature across different timeframes – pre-2000, 2000-2010, and post-2010. In Fig. 1, we show the distribution of sources across different time frames. Paying attention to references prior to 2000 is particularly important because they include some well-established data quality metrics that remain relevant today. These early efforts provide the context and insights into an evolution of data readiness metrics over the years. Before 2000, work primarily focused on general data quality considerations without specific emphasis on applications in AI. From 2000 to 2010, AI research was focused on foundational computational methods and concepts. Post-2010, with the rise of big data and machine learning technologies, a notable shift toward creating metrics to assess data preparedness emerged, specifically for AI applications.
2.1. Existing Surveys and Gaps
Existing surveys on data quality ((Batini and Scannapieco, 2006; Elmagarmid et al., 2007; Sidi et al., 2012; Thung and Raveendran, 2009; Lin and Jay Kuo, 2011)) primarily focused on traditional dimensions like completeness, correctness, timeliness, and also the quality of textual and multimedia data, providing a strong foundation for understanding the general challenges in the field. However, with the rise of AI, there has been an increasing emphasis on AI-specific concerns like feature relevance, class imbalance, mislabeled data, privacy, and fairness ((Wagner and Eckhoff, 2018; Ntoutsi et al., 2020; Li et al., 2017; Priestley et al., 2023; Shahbazi et al., 2023)). These emerging factors are critical alongside conventional quality benchmarks. Recognizing the importance of both aspects, our survey aims to address this shift towards a comprehensive view of data readiness, where traditional and AI-focused dimensions are equally essential.
@@ 2.1. Existing Surveys and Gaps
Toward gaining an understanding of metrics for data readiness, a few studies surveyed data pre-processing stages in preparing data for AI. Priestley et al. (2023) conducted a study highlighting the role of decision-makers and practitioners in improving data-focused practices. They highlighted the significance of data cleaning and pre-processing stages, including feature selection, duplicate elimination, outlier removal, consistency assurance, and handling of missing values. Documentation of these pre-processing steps was essential to ensure compatibility and identify potential dependencies and information leakages among features. Their insights, derived from an extensive literature survey, laid the foundation for recognizing the pivotal role of data readiness in AI applications. Yalaoui and Boukhedouma (2021) presented a survey paper on data quality evaluation and enhancement. They compared existing frameworks, models, and methods for data quality evaluation and enhancement, and identified challenges.
Batini and Scannapieco’s (Batini and Scannapieco, 2006) book offers a comprehensive introduction to a broad range of data quality issues. This book provides a state of the art overview of data quality measurement practices by using probability theory, data mining, statistical data analysis, and machine learning.
@@
A few studies have looked into metrics for unstructured data. Elmagarmid et al. (2007) explored similarity metrics for duplicate record detection, specifically focusing on challenges posed by typographical variations in string data. Li et al. (2017) concentrated on feature selection metrics, categorizing traditional approaches into wrapper, filter, embedded, and hybrid methods. Furthermore, Forman (2003) explored feature selection metrics for text classification, examining their performance. The studies on image quality measures (Thung and Raveendran, 2009) and perceptual visual quality metrics (Lin and Jay Kuo, 2011) are important in understanding DRAI for unstructured data, particularly in visual data, where assessment of image quality and perceptual metrics are essential for effective AI applications.
Figure 2. 360° View of Mapping Data Readiness Dimensions for AI
@@
In the context of bias and fairness, Shahbazi et al. (2023) provided a survey of techniques focused on identifying and mitigating representation bias across diverse data types, such as structured data, image data, textual data, and graph data. They defined the problem and discussed causes and methods for measuring and quantifying representation bias in structured and unstructured datasets. Addressing the issue of discrimination in DRAI systems, Ntoutsi et al. (2020) focused on the challenges and implications of biased data on the fairness and accuracy of AI-based decision-making. They emphasized the importance of understanding and mitigating biases in data to prevent the continuation of discriminatory practices. Similarly, in the context of privacy, Wagner and Eckhoff (2018) reviewed over $80$ privacy metrics across six domains, categorizing them based on the aspect of privacy measured, required inputs, and data types needing protection. They identified research gaps in areas such as metric combination and interdependent privacy, proposing a method for selecting appropriate metrics through nine key questions. Emphasizing the importance of employing multiple metrics to address diverse privacy aspects, the paper serves as a reference guide and toolbox for privacy researchers, aiding informed choices in metric selection for specific scenarios.
@@
Our study aims to contribute to the evolving field of DRAI metrics, specifically focusing on structured and unstructured data metrics. While previous works have explored specific dimensions or concentrated on individual metrics, we address the lack of a comprehensive study including numerous DRAI dimensions. There are broader perspectives in addressing DRAI applications. We identify numerous important factors that define data readiness, such as data preparation, privacy leakage evaluation, data discrimination evaluation, compliance to FAIR principles, mislabeled data detection, feature relevancy analysis, bias-related issues, and quality evaluation of speech and multimedia data. To survey existing literature in these dimensions and to identify gaps, our effort aims to advance the field of data readiness for AI (DRAI) and inform best practices for evaluating the suitability of data for AI applications. Given the growing importance of DRAI, our work fills a critical gap in the published literature and gives insights to practitioners and decision-makers in the field. As illustrated in Fig.  2, the 360° plot view of data readiness incorporates a range of metrics that reflect the ongoing discussion in the literature for both structured and unstructured data. Additionally, we also define DRAI by introducing key pillars and the DRAI metrics for each pillar.